# Making it to First: The Random Access Problem in DNA Storage

### v1

@@ 5 Conclusion and Future Research
In this paper, we addressed the random access problem in DNA storage. We determined the exact values of $T_{q}(2)$ and $T(2)$ and we extended the construction presented in [17]. Our analysis demonstrated that this generalization achieves the best-known results to date for $k=3$ and $k=4$. Despite these advancements, several intriguing questions remain open, which we aim to explore in future work. A couple of these directions are the following:
@@
    We have shown that the construction in Section 4 performs well for $k=3$ and $k=4$ and in particular the random access expectation of the construction for $k=4$ is smaller than $k=3$. Does this expectation continue to decrease as $k$ increases, and if so, what is its limit? Additionally, From [4], we know that for positive integers $a$ and $k$, it holds that $\frac{ak}{ak}\leq\frac{T(k)}{k}$, and we wonder whether it holds that $\frac{T(k+1)}{k+1}<\frac{T(k)}{k}$ for all $k$? If that is true, what is the limit of $\frac{T(k)}{k}$ as $k$ tends to infinity?
@@
    In all previous constructions of matrices achieving a low random access expectation, the number of columns of weight $m$ was the same for every set of $m$ coordinates. In other words, the marginal distribution of vectors of weight $m$ was uniform. However, it had not been formally proven that this uniformity is the optimal structure for minimizing $T_{\max}(G)$. For $k=2$, we have shown that the optimal matrix adheres to this uniform structure, but does this hold true for general $k$?
@@ Appendix A Appendix
For the reader’s convenience, we restate the statement of the theorem before presenting the corresponding proof.
See 3
@@
$\lim_{x\xrightarrow{}\infty}\frac{2}{3x(1+\alpha)-2}=0.$
@@
$\lim_{x\xrightarrow{}\infty}\frac{\alpha x-1}{3x(1+\alpha)-1}=\frac{\alpha}{3+3\alpha}.$
@@
$\lim_{x\xrightarrow{}\infty}\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}$
$\leq\lim_{x\xrightarrow{}\infty}\frac{x+2\alpha x}{3x(1+\alpha)-1}\cdot\frac{x+2\alpha x-1}{3x(1+\alpha)-2}(\frac{(\frac{1+2\alpha}{3+3\alpha})^{x+2\alpha x-1}-1}{\frac{-2-\alpha}{3+3\alpha}}-1)$
@@
Where $(a)$ holds because (${\alpha x\choose{s-1}}=\frac{s}{\alpha x+1}{\alpha x+1\choose s}$ and $(b)$ is due to
(${\alpha x+1\choose s}{3x(1+\alpha)-1\choose\alpha x+1}={3x(1+\alpha)-1-s\choose\alpha x+1-s}{3x(1+\alpha)-1\choose s}.$
@@
Now, by applying Pascal’s triangle recursively on (${3x+3\alpha x-3\choose\alpha x-2}$ we have that

### v2

@@ 5 Conclusion and Future Research
In this paper, we addressed the random access problem in DNA storage. We determined the exact values of $T_{q}(2)$ and $T(2)$ and we extended the construction presented in [19]. Our analysis demonstrated that this generalization achieves the best-known results to date. Despite these advancements, several intriguing questions remain open, which we aim to explore in future work. A few of these directions are the following:
@@
    From [4], we know that for positive integers $a$ and $k$, it holds that $\frac{T(ak)}{ak}\leq\frac{T(k)}{k}$, and we wonder whether it holds that $\frac{T(k+1)}{k+1}\leq\frac{T(k)}{k}$ for all $k$? If that is true, what is the limit of $\frac{T(k)}{k}$ as $k$ tends to infinity?
@@
    In all existing constructions that achieve a small random-access expectation, including those in the present work, the number of columns of weight $m$ is the same for every set of $m$ coordinates. In other words, the marginal distribution of vectors of weight $m$ is uniform. However, it had not been formally proven that this uniformity is the optimal structure for minimizing $T_{\max}(G)$. For $k=2$, we have shown that the optimal matrix adheres to this uniform structure, but is this true for general $k$? Even when we restrict ourselves to matrices whose columns have weight at most two (as we did in Construction 1), it is still unknown whether the symmetric pattern is truly best. We conjecture that it is, and establishing this rigorously is an important direction for future work.
  3. 3.
    The best lower bounds on $T_{q}(n,k)$ to date were obtained in the seminal work [4]: $T_{q}(n,k)\geq n-\frac{n(n-k)}{n}(H_{n}-H_{n-k})$ and $T_{q}(n,k)\geq\frac{k+1}{2}$. Despite their significance, these bounds remain well below the best upper bounds currently known. Sharpening them—or closing the gap from the upper-bound side—constitutes an important avenue for future research.
  4. 4.
    In this work, we analyzed only the substitution of $p=P$ in Theorem 3. A natural direction for future research is to optimize $p$ and $P$ given $k$, in order to minimize the expression in Theorem 3 and obtain a better upper bound on $T(k)$.
@@ Appendix A Appendix
For the reader’s convenience, we restate the statement of the lemma before presenting the corresponding proof.
See 5
@@
$\lim_{x\to{}\infty}\frac{2}{3x(1+\alpha)-2}=0.$
@@
$\lim_{x\to{}\infty}\frac{\alpha x-1}{3x(1+\alpha)-1}=\frac{\alpha}{3+3\alpha}.$
@@
$\lim_{x\to{}\infty}\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}$
$\leq\lim_{x\to{}\infty}\frac{x+2\alpha x}{3x(1+\alpha)-1}\cdot\frac{x+2\alpha x-1}{3x(1+\alpha)-2}(\frac{(\frac{1+2\alpha}{3+3\alpha})^{x+2\alpha x-1}-1}{\frac{-2-\alpha}{3+3\alpha}}-1)$
@@
Where $(a)$ holds because ${\alpha x\choose{s-1}}=\frac{s}{\alpha x+1}{\alpha x+1\choose s}$ and $(b)$ is due to
${\alpha x+1\choose s}{3x(1+\alpha)-1\choose\alpha x+1}={3x(1+\alpha)-1-s\choose\alpha x+1-s}{3x(1+\alpha)-1\choose s}$.
@@
Now, by applying Pascal’s triangle recursively on $3x+3\alpha x-3\choose\alpha x-2$} we have that